OpenStudy (anonymous):

Sarah earned some money taking care of a dog. She charged $8.50 for washing the dog and $7 for walking the dog. She spent $1.15 on shampoo to wash the dog. What was Sarah’s profit from taking care of the dog?

OpenStudy (anonymous):

yes, just add all the numbers up (: what are the answer choices ?

OpenStudy (anonymous):

nuthing it is a box with a $

OpenStudy (ilovecake):

I think the answer would be $13.35. I said that on your last post. I guess you were not there.

OpenStudy (anonymous):

thanks

OpenStudy (ilovecake):

yw

OpenStudy (anonymous):

one more

OpenStudy (ilovecake):

ok

OpenStudy (anonymous):

Nolan lost 12.5 points in the final round of a game, bringing his score down to –8. Suppose p represents Nolan’s points before he lost the final round. Which equation can be used to represent this situation? A. p – (–8) = –12.5 B. p + 12.5 = –8 C. p – 12.5 = –8 D. p + 8 = 12.5

OpenStudy (ilovecake):

?

OpenStudy (anonymous):

o sorry hold on

OpenStudy (anonymous):

Nolan lost 12.5 points in the final round of a game, bringing his score down to –8. Suppose p represents Nolan’s points before he lost the final round. Which equation can be used to represent this situation?

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

Nolan lost 12.5 points in the final round of a game, bringing his score down to –8. Suppose p represents Nolan’s points before he lost the final round.

OpenStudy (anonymous):

o do not pay a tenshion to the marks

OpenStudy (ilovecake):

p – 12.5 = –8